OpenStudy (anonymous):

Define a function f : R → R by the formula f(x) = 4x − 8. Prove that f is one-to-one. Let a, b is in R, and suppose that f(a) = f(b). Then 4a − 8 =

OpenStudy (amistre64):

it pretty much writes itsefl ...

OpenStudy (zzr0ck3r):

assume f(a) = f(b) then 4a-8 = 4b-8 4a=4b a=b QED